OpenStudy (isaacsgames):

Solve for z z/12 = 6 3/4

OpenStudy (isaacsgames):

@Nnesha @seb.cal @pooja195 @porkbuns @iambatman @texaschic101 @suppa_lova

OpenStudy (isaacsgames):

Fan and Meadl

OpenStudy (isaacsgames):

I have to go soon!

OpenStudy (isaacsgames):

@calculusxy

OpenStudy (isaacsgames):

@Hero

OpenStudy (isaacsgames):

@Owlcoffee

OpenStudy (isaacsgames):

@mathwiz916

OpenStudy (anonymous):

\[\frac{z}{6}=6*\frac{3}{4}\] \[\frac{z}{6}=\frac{18}{4}\] multiply by 6: \[z=\frac{108}{4}=27\]

OpenStudy (isaacsgames):

Still confused

OpenStudy (isaacsgames):

OH, now I get it

OpenStudy (isaacsgames):

Thanks!

OpenStudy (anonymous):

You're welcome